Ask
your own question, for FREE!
Mathematics
7 Online
https://pacyber-li.brainhoney.com/Resource/34857366,0/Assets/candy.jpg Kareem closes his eyes and picks a candy at random from the group above and eats it. Later he goes back to choose another candy at random. What is the chance of him getting a yellow then a purple? P(yellow) * P(purple after yellow) Write your answer as a fraction in simplest form. 1/5 1/9 1/20
Still Need Help?
Join the QuestionCove community and study together with friends!
pa cyber homie hey
i saw pa cyber . brainhoney and i was like heeey. I don't know the answer but what grade ar eyou in?
Kareem has a 1/5 chance of choosing a yellow candy. After he removes a candy, he has a 1/4 chance of getting a purple candy. To find the probability of these events, we multiply the two together. 1/4*1/5=1/20
9th ;p
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
27 minutes ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal